# Mimicking Better by Matching the Approximate Action Distribution

### v2

@@ 2 Related Work
Within the ILD setting a number of works take an alternative approach to address the non-availability of actions; instead of operating over state-state transitions they seek to infer the non-available actions. They do so by learning an inverse dynamics model, use it to infer the actions and continue with the application of standard ILD methods (Torabi et al., 2018; Hanna & Stone, 2017; Nair et al., 2017; Pavse et al., 2019; Al-Hafez et al., 2023). The availability of an inverse dynamics model that can account for the non-available actions opens the door for bringing the sample efficiency properties of behavioral cloning within the ILO context. (Zhu et al., 2021) follow such a path combining an off-policy imitation learning algorithm (Kostrikov et al., 2018) with a BC-like regularizer and demonstrate sample efficiency gains. Similarly, Hybrid-RL (Guo et al., 2019) merges standard RL with a BC-like regularizer; however, it assumes access to the expert’s reward function, which limits its broader applicability.
@@ 4.2 Controlling the Policy Learning with the Inverse Dynamics Model
MAAD-AIL |
MAAD-TM |
MAAD-OT |
@@ 5 Experiments
To evaluate MAAD, we have conducted experiments on complex control tasks from the MuJoCo suite of environments (Todorov et al., 2012). We have used 6 locomotion tasks from DeepMind Control Suite and 5 from OpenAI Gym. The two suites differ on various aspects, such as their initial state distributions and termination criteria, posing their specific challenges to the learning algorithms, more details in Section B.1. We collected expert trajectories from a policy trained using PPO (Schulman et al., 2017) on each MuJoCo task. Then we used the collected trajectories to train several imitation learning baseline models and compare them against different flavors of our model. In the baseline models, we included not only ILO methods but also some ILD ones which have access to expert actions, these later ones should be seen as the best possible achievable performance if actions were observed. More details about environments, hyperparameters and training data and implementation can be found in Section B in appendix, and our code is openly available:

### v3

@@ 2 Related Work
Within the ILO setting a number of works take an alternative approach to address the non-availability of actions; instead of operating over state-state transitions they seek to infer the non-available actions. They do so by learning an inverse dynamics model, use it to infer the actions and continue with the application of standard ILD methods (Torabi et al., 2018; Hanna & Stone, 2017; Nair et al., 2017; Pavse et al., 2019; Al-Hafez et al., 2023; Radosavovic et al., 2020). Additionally, it can also be used to pretrain a policy (Brandfonbrener et al., 2023). The availability of an inverse dynamics model that can account for the non-available actions opens the door for bringing the sample efficiency properties of behavioral cloning within the ILO context. (Zhu et al., 2021) follow such a path combining an off-policy imitation learning algorithm (Kostrikov et al., 2018) with a BC-like regularizer and demonstrate sample efficiency gains. Similarly, Hybrid-RL (Guo et al., 2019) merges standard RL with a BC-like regularizer; however, it assumes access to the expert’s reward function, which limits its broader applicability.
@@ 4.2 Controlling the Policy Learning with the Inverse Dynamics Model
| MAAD-AIL | MAAD-TM | MAAD-OT |
@@ 5 Experiments
To evaluate MAAD, we have conducted experiments on complex control tasks from the MuJoCo suite of environments (Todorov et al., 2012). We have used 6 locomotion tasks from DeepMind Control Suite and 5 from OpenAI Gym. The two suites differ on various aspects, such as their initial state distributions and termination criteria, posing their specific challenges to the learning algorithms, more details in Section B.1. We collected expert trajectories from a policy trained using PPO (Schulman et al., 2017) on each MuJoCo task. Then we used the collected trajectories to train several imitation learning baseline models and compare them against different flavors of our model. In the baseline models, we included not only ILO methods but also some ILD ones which have access to expert actions, these later ones should be seen as the best possible achievable performance if actions were observed. More details about environments, hyperparameters and training data and implementation can be found in Section B in appendix, and our code is openly available:https://github.com/jacr13/MAAD.
@@ 6 Conclusion
Acknowledgements
This work was partially supported by the Swiss National Science Foundation grant number CSSII5_177179 “Modeling pathological gait resulting from motor impairment”.
Impact Statement
This paper presents work whose goal is to advance the field of Reinforcement Learning, specifically in the domain of imitation learning from observations, by introducing a regularizer that encourages the agent to reproduce physically-plausible expert actions (estimated via a learned inverse dynamics model). Our research was conducted using simulators and expert data collected from a trained agents, with no real-world data involved. While our work may have various societal implications, we do not believe any particular consequences need to be highlighted here.